1) For a fair six-sided die, determine the following probabilites: A. Getting a 2 and a 6 on two independent die rolls. B. NOT getting a 4 on a single die roll.
Answer #1

P(getting a two and getting a six) = P( getting a 2 first and then 6) + P( getting a 6 first and then 6)

= 2* P(getting a 2) * P(getting a 6) = 2*( 1/6) * (1/6) = 1/18 = 0.056

P (not getting a 4 ) = 1 - P(getting a 4) = 1 - 1/6 =5/6 =0.833